# Self-distributive structures in physics

### v2

@@ Abstract.
It is an important feature of our existing physical theories that observables generate one-parameter groups of transformations. In classical Hamiltonian mechanics and quantum mechanics, this is due to the fact that the observables form a Lie algebra, and it manifests itself in Noether’s theorem. In this paper, we introduce Lie quandles as the minimal mathematical structure needed to express the idea that observables generate transformations. This is based on the notion of a quandle used most famously in knot theory, whose main defining property is the self-distributivity equation $x\triangleright(y\triangleright z)=(x\triangleright y)\triangleright(x\triangleright z)$. We argue that Lie quandles can be thought of as nonlinear generalizations of Lie algebras.
@@
Acknowledgements. We thank Benjamin Walder for helpful discussions on Lie quandles and Markus Szymik for valuable pointers to the literature.
@@ 1. Introduction
Observables play a dual role in our existing physical theories [1]. On the one hand, they are the quantities that we think of as being measurable in principle. On the other hand, they are also exactly those quantities that generate groups of transformations, including symmetries and the dynamics of a theory. This surprising dual role is especially apparent and well-understood in Hamiltonian mechanics, where it is closely tied to the symplectic structure of phase space, and in quantum mechanics, where it manifests itself in the correspondence between self-adjoint operators and one-parameter groups of unitaries (Stone’s theorem). In both cases, this duality is what underlies Noether’s theorem: an observable is a conserved quantity if and only if the transformation group that it generates is a group of symmetries.
This fact that observables generate transformations is not an a priori necessary feature of physical theories. In fact, it is not difficult to come up with candidate physical theories which violate it. For example, this is generically the case for the general probabilistic theories studied in quantum foundations [2, 3]: in a general probabilistic theory, a state space is given by a convex body in $\mathbb{R}^{n}$, and a generic such object has no symmetries at all, and in particular no nontrivial one-parameter groups of transformation. Another class of theories in which observables cannot generate transformations in the way that we are used to are theories with non-reversible time evolution. What we are used to is that if dynamics is generated by an observable denoted $H$, then we would expect there to be another observable $-H$ which generates the time reversed dynamics.
We are thus led to conclude that having a well-behaved correspondence between observables and one-parameter groups of transformations is a distinguishing feature of those theories that describe the physics of our universe (as far as we understand it). Suppose now that we want to develop other theories—perhaps more fundamental ones—that follow this very principle. Then which mathematical structure do we need in order to capture this feature? In Hamiltonian mechanics and quantum mechanics, this principle is encoded in the fact that the spaces of observables is a Lie algebra. This leads to the desired feature because any observable $H$ generates a one-parameter group of transformations on the space of observables itself determined by the equation
@@
This intriguingly self-referential equation is a form of self-distributivity, which famously appears in the theory of self-distributive structures like quandles [4, 5]. Surprisingly, the role of such self-distributivity in physics does not seem to have been appreciated before [^1] One notable exception is a blog post by Daniel Moskovich [6], which expresses the idea that associative algebraic structures are geometric in nature, while self-distributive operations are appropriate for ‘information physics’. However, he does not seem to have recognized the role of self-distributivity in the correspondence between observables and transformation groups., and quandles have hitherto mainly been considered in pure mathematics [^2] Their most well-known application is to knot theory [7, 4, 8]. However, the earliest appearance of (involutive) quandles seems to be as algebraic structures abstracting reflections [9], and they have also been used in number theory [10] and real algebra [11]. Brieskorn’s early survey [12] discusses relations to braid groups, root systems and singularity theory.. Hence the main purpose of this paper is to point out the relevance of self-distributivity in physics.
There is another appearance of self-distributivity in physics which underlines its relevance, but now at the level of states rather than observables. In quantum theory, forming convex combinations or mixtures of (mixed) states has been recognized as a crucial piece of structure already by von Neumann [13, Ch. IV], and has since become central through the framework of general probabilistic theories. Now if we consider convex combinations operations in the form
$x\ast_{t}y\coloneqq(1-e^{-t})x+e^{t}y,$
@@
which is an equation that generally does not hold for the $\triangleright_{t}$ operations.
@@
While our notion of Lie quandle and its physical interpretation in terms of observables generating transformations seem to be new, the importance of self-distributivity in physics has been recognized in a blog post by Moskovich [6]. There also are a number of existing works relating self-distributive structures with Lie theory:
@@
    The thesis of Crans [14, Section 3.2.4] considers unital spindles and quandles internal to the category of coalgebras, and uses these structures as an intermediate notion between Lie groups and Lie algebras. Since the underlying set of her quandles correspond to the symmetric power of a Lie algebra, this seems quite different from our Lie quandles (of which the simplest examples are Lie algebras themselves).
@@
    Smooth quandles were introduced by Ishikawa [15] [^3] There is a more general notion of quandles in categories with finite products [16], which yields smooth quandles upon taking the category to be the category of manifolds.; we also refer to the thesis of Yonemura [17] for a detailed exposition. Smooth quandles are quandles internal to the category of manifolds (meaning that the underlying set is a manifold and the quandle operation is smooth). This is more closely related to Lie quandles: a Lie quandle becomes a smooth quandle if one forgets all operations $\triangleright_{t}$ for $t\neq 1$.
@@ 2. Self-distributivity
To begin a more detailed discussion, let us first consider the mathematics of self-distributivity in its essence. We will not do anything original here and we refer to literature like [18, 4, 14, 5] for more details.
Definition 2.1 (e.g. [18, 14]).
A shelf is a set $S$ together with a binary operation normal-$\triangleright$ such that
@@
Here we have adopted the terminology of Crans [14, Definition 52] [^4] We the attribute ‘left’ since we will not consider right shelves.. A good way to think about this equation is as saying that the map $x\triangleright-:S\to S$ is a homomorphism with respect to $\triangleright$ itself for every $x\in S$. In this way, shelves axiomatize the idea of a set acting on ‘itshelf’.
@@
may be the historically first example of a shelf [18, p. 481]. Note that every element acts by permuting the other two.
@@
    For another example, take $S\coloneqq\mathbb{R}^{2}$ and define $x\triangleright y$ to be the point obtained by reflecting $y$ across $x$. This satisfies self-distributivity due to the fact that any Euclidean transformation preserves distances, the point $x\triangleright y$ can be characterized uniquely in terms of its distances from $x$ and $y$ [^5] Namely as the unique point whose distance from $x$ equals the distance of $x$ and $y$, and whose distance from $y$ is twice that., and reflection at $x$ is a Euclidean transformation.
@@
    More generally, for any fixed $\theta\in[0,2\pi)$ by obtain a spindle by defining $x\triangleright y$ to be the point obtained by rotating $y$ around $x$ by $\theta$; taking $\theta=\pi$ then recovers the previous example.
@@
    Similarly, we can work with points on the sphere $S^{2}$ and define $x\triangleright y$ to be the point obtained by rotating $y$ around $x$ by $\theta$ [19, Definition 4.3].
@@
    We can also take $S$ to be any symmetric space, such as the sphere $S^{2}$, and define $x\triangleright y$ to be the point obtained by reflecting $y$ across $x$ [5, Example 56] [^6] Note that [5] considers right self-distributivity rather than the left version of Definition 2.1, so that their $x\triangleright y$ corresponds to our $y\triangleright x$..
@@
The following example is a simple variation on the standard Alexander quandle from the theory of knot invariants [5, Example 67]. It also appears in Moskovich’s blog post [6] and has probably been known before.
@@
Moreover, it can be shown that $\triangleright_{-1}$ equips the original set $Q$ with another quandle structure, and $\triangleright_{1}$ and $\triangleright_{-1}$ also mutually distribute over each other [5, p. 95]. These statements are observations that naturally lead to the Lie quandles of Section 3 if one considers an entire family of operations $\triangleright_{t}$ indexed by $t\in\mathbb{R}$.
@@
    Assuming $s\in(0,1)$, a convex set $C\subseteq\mathbb{R}^{n}$ is typically not a quandle with respect to (2.3), since for given $x,z\in C$ the point $y$ with $(1-s)x+sy=z$ need not belong to $C$ again. In fact, it can be show that $C$ is a quandle if and only if it is an affine subspace of $\mathbb{R}^{n}$.
@@
    This construction works the same way if $G$ is replaced by any quandle and the action on $X$ by a quandle action in the sense of [20, § 3.2].
Hitherto, the theory of quandles has been primarily applied to knot theory. We refer to the original papers [7, 4, 8] as well as to the more recent textbook [5] for expositions of this application. [^7] A reader who knows about anyon physics may wonder whether anyons provide another way in which quandles appear in physics. As far as we can see this is not the case, since anyons are modelled by braided fusion categories, and these do not seem to bear any direct relation (with physical significance) to quandles.
@@ 3. Lie quandles
As indicated in the introduction, we now introduce a Lie-theoretic version of quandles. From our point of view, these are the minimal structures needed to formalize the idea that observables of a physical theory form a set such that every elements acts on the set itself as a one-parameter transformation group.
@@ A Lie quandle is a smooth manifold $Q$ together with an operation
for any matrix $Y$ and any invertible $A$ [21, Proposition 2.3(6)]. We apply this with $A=e^{isX}$ in the third step of the following sequence of equations:
@@
The characterization (3.4) of the Lie quandle structure implies that every (real) Lie algebra $\mathfrak{g}\subseteq M_{n}(\mathbb{C})$ is a Lie quandle with respect to the operations $\triangleright_{t}$ restricted from all matrices to $\mathfrak{g}$. By Ado’s theorem, we can embed every Lie algebra into $M_{n}(\mathbb{C})$ for some $n$, and therefore every Lie algebra is a Lie quandle in a canonical way. If $G$ is any Lie group with Lie algebra $\mathfrak{g}$, then this Lie quandle structure on $\mathfrak{g}$ can also be written in terms of the exponential map $\exp:\mathfrak{g}\to G$ and the adjoint action of $G$ on $\mathfrak{g}$ [21, §4.3.3] as
@@
We introduce a Lie-theoretic version of the sphere quandle appearing in Example 2.4Item (ii) by considering all angles of rotation at the same time. This amounts to taking $Q\coloneqq S^{2}$, and for given points $x,y\in S^{2}$, we define $x\triangleright_{t}y$ as the point obtained by rotating $y$ by an angle $t$ around the axis having $x\in S^{2}$ as its north pole. [^9] Whether one rotates in the positive or negative direction is then irrelevant as long as it is done consistently.
@@
The correspondence between observables and one-parameter groups of transformations plays a crucial role in Noether’s theorem. [^10] We refer to [22, Theorem 5.58] for a general version of Noether’s theorem in classical field theory. We can now understand this result by phrasing a formulation of Noether’s theorem due to Baez [1] in the framework of Lie quandles.
@@ 4. Conclusion
We have tried to argue that self-distributive structures deserve to play a prominent role in physics, upending the usual focus on associative structures. This is in line with Moskovich’s idea that associative algebraic structures are geometric in nature, while self-distributive operations are appropriate for ‘information physics’ [6]. The basic observation is that observables form a self-distributive structure with respect to every observable generating a one-parameter transformation group on the space of observables, and we have proposed a definition of Lie quandle to formalize this structure more generally, and we have argued that Lie quandles are nonlinear generalizations of Lie algebras.

### v3

# quandle_physics.bib

@@ Abstract.
It is an important feature of our existing physical theories that observables generate one-parameter groups of transformations. In classical Hamiltonian mechanics and quantum mechanics, this is due to the fact that the observables form a Lie algebra, and it manifests itself in Noether’s theorem. In this paper, we propose Lie quandles as the minimal mathematical structure needed to express the idea that observables generate transformations. This is based on the notion of a quandle used most famously in knot theory, whose main defining property is the self-distributivity equation $x\triangleright(y\triangleright z)=(x\triangleright y)\triangleright(x\triangleright z)$. We argue that Lie quandles can be thought of as nonlinear generalizations of Lie algebras.
@@
Acknowledgements. We thank Benjamin Walder for helpful discussions on Lie quandles, John Baez and Markus Szymik for further discussion and valuable pointers to the literature and the anonymous referee for suggesting interesting examples of Noether quandles.
@@ 1. Introduction
Observables play a dual role in our existing physical theories [baez]. On the one hand, they are the quantities that we think of as being measurable in principle. On the other hand, they are also exactly those quantities that generate groups of transformations, including symmetries and the dynamics of a theory. This surprising dual role is especially apparent and well-understood in Hamiltonian mechanics, where it is closely tied to the symplectic structure of phase space, and in quantum mechanics, where it manifests itself in the correspondence between self-adjoint operators and one-parameter groups of unitaries (Stone’s theorem). In both cases, this duality is what underlies Noether’s theorem: an observable is a conserved quantity if and only if the transformation group that it generates is a group of symmetries.
This fact that observables generate transformations is not an a priori necessary feature of physical theories. In fact, it is not difficult to come up with candidate physical theories which violate it. For example, this is generically the case for the general probabilistic theories studied in quantum foundations [barrett, plavala]: in a general probabilistic theory, a state space is given by a convex body in $\mathbb{R}^{n}$, and a generic such object has no symmetries at all, and in particular no nontrivial one-parameter groups of transformation. Another class of theories in which observables cannot generate transformations in the way that we are used to are theories with non-reversible time evolution. What we are used to is that if dynamics is generated by an observable denoted $H$, then we would expect there to be another observable $-H$ which generates the time reversed dynamics.
We are thus led to conclude that having a well-behaved correspondence between observables and one-parameter groups of transformations is a distinguishing feature of those theories that describe the physics of our universe (as far as we understand it). Suppose now that we want to develop other theories—perhaps more fundamental ones—that follow this very principle. Then which mathematical structure do we need in order to capture this feature? In Hamiltonian mechanics and quantum mechanics, this principle is encoded in the fact that the spaces of observables is a Lie algebra. This leads to the desired feature because any observable $H$ generates a one-parameter group of transformations on the very space of observables itself determined by the equation
@@
This intriguingly self-referential equation is a form of self-distributivity, which famously appears in the theory of self-distributive structures like quandles [joyce, EN]. Surprisingly, the role of such self-distributivity in physics does not seem to have been appreciated before [^1] One notable exception is a blog post by Daniel Moskovich [moskovich], which expresses the idea that associative algebraic structures are geometric in nature, while self-distributive operations are appropriate for ‘information physics’. However, he does not seem to have recognized the role of self-distributivity in the correspondence between observables and transformation groups., and quandles have hitherto mainly been considered in pure mathematics [^2] Their most well-known application is to knot theory [burde, joyce, matveev]. However, the earliest appearance of (involutive) quandles seems to be as algebraic structures abstracting reflections [takasaki], and they have also been used in number theory [LL] and real algebra [szymik]. Brieskorn’s early survey [brieskorn] discusses relations to braid groups, root systems and singularity theory.. Hence the main purpose of this paper is to point out the relevance of self-distributivity in physics.
There is another appearance of self-distributivity in physics which underlines its relevance, but now at the level of states rather than observables. In quantum theory, forming convex combinations or mixtures of (mixed) states has been recognized as a crucial piece of structure already by von Neumann [vonneumann, Ch. IV], and has since become central through the framework of general probabilistic theories. Now if we consider convex combinations operations in the form
$x\ast_{t}y\coloneqq(1-e^{-t})x+e^{-t}y,$
@@
which is an equation that typically does not hold for the $\triangleright_{t}$ operations.
@@
While our notion of Lie quandle and its physical interpretation in terms of observables generating transformations seem to be new, the importance of self-distributivity in physics has been recognized in a blog post by Moskovich [moskovich]. There also are a number of existing works relating self-distributive structures with Lie theory:
@@
    The thesis of Crans [crans, Section 3.2.4] considers unital spindles and quandles internal to the category of coalgebras, and uses these structures as an intermediate notion between Lie groups and Lie algebras. Since the underlying set of Crans’ quandles correspond to the symmetric powers of Lie algebras, this seems quite different from our Lie quandles (of which the simplest examples are Lie algebras themselves).
@@
    Smooth quandles were introduced by Ishikawa [ishikawa] [^3] There is a more general notion of quandles in categories with finite products [grosfjeld], which yields smooth quandles upon taking the category to be the category of manifolds.; we also refer to the thesis of Yonemura [yonemura] for a detailed exposition. Smooth quandles are quandles internal to the category of manifolds (meaning that the underlying set is a manifold and the quandle operation is smooth). This is more closely related to Lie quandles: a Lie quandle becomes a smooth quandle if one forgets all operations $\triangleright_{t}$ for $t\neq 1$.
@@ 2. Self-distributivity
To begin a more detailed discussion, let us first consider the mathematics of self-distributivity in its essence. We will not do anything original here, and we refer to literature like [dehornoy, joyce, crans, EN] for more details.
Definition 2.1 (e.g. [dehornoy, crans]).
A shelf is a set $S$ together with a binary operation $\triangleright$ such that
@@
Here we have adopted the terminology of Crans [crans, Definition 52] [^4] We omit the attribute ‘left’ since we will not consider right shelves.. A good way to think about this equation is as saying that the map $x\triangleright-:S\to S$ is a homomorphism with respect to $\triangleright$ itself for every $x\in S$. In this way, shelves axiomatize the idea of a set acting on ‘itshelf’.
@@
may be the historically first example of a shelf [dehornoy, p. 481]. Note that every element acts by permuting the other two.
@@
    For another example, take $S\coloneqq\mathbb{R}^{2}$ and define $x\triangleright y$ to be the point obtained by point reflecting $y$ across $x$. This satisfies self-distributivity due to the fact that any Euclidean transformation preserves distances, the point $x\triangleright y$ can be characterized uniquely in terms of its distances from $x$ and $y$, [^5] Namely as the unique point whose distance from $x$ equals the distance of $x$ and $y$, and whose distance from $y$ is twice that. and reflection at $x$ is a Euclidean transformation.
@@
    More generally, we can take $S$ to be any symmetric space, such as the sphere $S^{2}$, and define $x\triangleright y$ to be the point obtained by reflecting $y$ across $x$ [EN, Example 56] [^6] Note that [EN] considers right self-distributivity rather than the left version of Definition 2.1, so that their $x\triangleright y$ corresponds to our $y\triangleright x$..
@@
    For any fixed $\theta\in[0,2\pi)$, we get another spindle structure on $\mathbb{R}^{2}$ by defining $x\triangleright y$ to be the point obtained by rotating $y$ around $x$ by $\theta$; taking $\theta=\pi$ then recovers the previous example.
@@
    Similarly, we can work with points on the sphere $S^{2}$ and define $x\triangleright y$ to be the point obtained by rotating $y$ around $x$ by $\theta$ [CS, Definition 4.3].
@@
The following example is a simple variation on the standard Alexander quandle from the theory of knot invariants [EN, Example 67]. It also appears in Moskovich’s blog post [moskovich] and has probably been known before.
@@
Moreover, it is easily seen that $\triangleright_{-1}$ equips the original set $Q$ with another quandle structure, [^7] Some authors use the opposite notational convention by introducing an operation $\triangleleft$ given by $x\triangleleft y\coloneqq y\triangleright_{-1}x$, which then satisfies the right self-distributivity equation $(x\triangleleft y)\triangleleft z=(x\triangleleft z)\triangleleft(y\triangleleft z)$ [crans, Section 3.1.1]. and $\triangleright_{1}$ and $\triangleright_{-1}$ also mutually distribute over each other [EN, p. 95]. These statements are observations that naturally lead to the Lie quandles of Section 3 if one considers an entire family of operations $\triangleright_{t}$ indexed by $t\in\mathbb{R}$.
@@
    Assuming $s\in(0,1)$, a convex set $C\subseteq\mathbb{R}^{n}$ is typically not a quandle with respect to (2.3), since for given $x,z\in C$ the point $y$ with $(1-s)x+sy=z$ need not belong to $C$ again. In fact, it is easy to see that such $C$ is a quandle if and only if it is an affine subspace of $\mathbb{R}^{n}$.
@@
    This construction works the same way if $G$ is replaced by any quandle and the action on $X$ by a quandle action in the sense of [elhamdadi, § 3.2].
Hitherto, the theory of quandles has been primarily applied to knot theory. We refer to the original papers [burde, joyce, matveev] as well as to the more recent textbook [EN] for expositions of this application. [^8] A reader who knows about anyon physics may wonder whether anyons provide another way in which quandles appear in physics. As far as we can see this is not the case, since anyons are modelled by braided fusion categories, and these do not seem to bear any direct relation (with physical significance) to quandles.
@@ 3. Lie quandles
As indicated in the introduction, we now introduce a Lie-theoretic version of quandles. From our point of view, these are the minimal structures needed to formalize the idea that observables of a physical theory form a set such that every elements acts on the set itself as a one-parameter group of transformations.
@@ A Lie quandle is a smooth manifold $Q$ together with an operation
for any matrix $Y$ and any invertible $A$ [hall, Proposition 2.3(6)]. We apply this with $A=e^{isX}$ in the third step of the following sequence of equations:
@@
The characterization (3.4) of the Lie quandle structure implies that every (real) Lie algebra $\mathfrak{g}\subseteq M_{n}(\mathbb{C})$ is a Lie quandle with respect to the operations $\triangleright_{t}$ restricted from all matrices to $\mathfrak{g}$. By Ado’s theorem, we can embed every Lie algebra into $M_{n}(\mathbb{C})$ for some $n$, and therefore every Lie algebra is a Lie quandle in a canonical way. If $G$ is any Lie group with Lie algebra $\mathfrak{g}$, then this Lie quandle structure on $\mathfrak{g}$ can also be written in terms of the exponential map $\exp:\mathfrak{g}\to G$ and the adjoint action of $G$ on $\mathfrak{g}$ [hall, §4.3.3] as
@@
We introduce the Bloch quandle, which is a Lie-theoretic version of the sphere quandle appearing in Example 2.4Item (ii) by considering all angles of rotation at the same time. This amounts to taking $Q\coloneqq S^{2}$, and for given points $x,y\in S^{2}$, we define $x\triangleright_{t}y$ as the point obtained by rotating $y$ by an angle $t$ around the axis having $x\in S^{2}$ as its north pole. [^10] Whether one rotates in the positive or negative direction is then irrelevant as long as it is done consistently.
@@
The correspondence between observables and one-parameter groups of transformations plays a crucial role in Noether’s theorem. [^11] We refer to [olver, Theorem 5.58] for a general version of Noether’s theorem in classical field theory. We can now understand this result by phrasing a formulation of Noether’s theorem due to Baez [baez] in the framework of Lie quandles.
@@
Example 3.6.
The Bloch quandle from Example 3.3 is a Noether quandle. One way to see this is to recall that it is a Lie subquandle of a Noether quandle, and the Noether property (3.6) is preserved by the passage to a Lie subquandle.
The analogously defined Lie quandles based on the Euclidean plane or the hyperbolic plane are Noether quandles as well. Indeed in all three cases, the equation $x\triangleright_{t}y=y$ for all $t\in\mathbb{R}$ is easily seen to be equivalent to $x=y$.
We owe the next example of a Noether quandles to the referee.
Example 3.7.
In a similar spirit to the Bloch quandle, fix distinct numbers $\lambda_{1},\ldots,\lambda_{n}\in\mathbb{R}$, and consider the set of all hermitian matrices in $M_{n}(\mathbb{C})$ with spectrum equal to $\{\lambda_{1},\ldots,\lambda_{n}\}$. By the spectral theorem, these matrices are in bijection with orthonormal bases of $\mathbb{C}^{n}$. Therefore this set is a submanifold of $M_{n}(\mathbb{C})$. Since conjugating by a unitary leaves the spectrum invariant, we obtain a Lie subquandle of $M_{n}(\mathbb{C})$, which is therefore a Noether quandle as well.
@@ 4. Conclusion
We have argued that self-distributive structures deserve to play a prominent role in physics, upending the usual focus on associative structures. This is in line with Moskovich’s idea that associative algebraic structures are geometric in nature, while self-distributive operations are appropriate for ‘information physics’ [moskovich]. The basic observation is that observables form a self-distributive structure with respect to every observable generating a one-parameter transformation group on the space of observables, and we have proposed a definition of Lie quandle to formalize this structure more generally, and we have argued that Lie quandles are nonlinear generalizations of Lie algebras.
@@
Funding and/or Conflicts of interests/Competing interests
No funding was received to assist with the preparation of this manuscript. The author has no competing interests to declare that are relevant to the content of this article.
[…]